# Boosting Adversarial Transferability for Hyperspectral Image Classification Using 3D Structure-invariant Transformation and Intermediate Feature Distance

### v1

@@ IV-D Ablation experiment
SpatialT SpectralT FDL Enlarge VGG-11 Resnet-18 VGG-19 Inc-V3 IncRes-V2
@@
SpatialT SpectralT FDL Enlarge VGG-11 Resnet-18 VGG-19 Inc-V3 IncRes-V2
@@
To gain insight into the performance of the method, we further analyze the influence of various hyperparameters on the experimental results. It mainly includes parameters such as block division, the number of copies, and the feature enlargement factor. The experiments were conducted mainly on the PaviaU dataset, with VGG-11 selected as the white-box model for generating adversarial examples, while the other models were treated as black-box models. The perturbation budget was set to 0.03.
The impact of the number of iterations on attack performance is shown in Fig. 5. The iterations range from 1 to 60. When the number of iterations exceeds 20 times, the improvement in the attack effect is limited. Therefore, to optimize computational costs while considering the attack’s efficacy, we set the number of iterations to 20.
The impact of the number of copies on attack performance is shown in Fig. 5. The number of replicated copies ranges from 1 to 20. As the number of transformed copies increases, the accuracy decreases most rapidly from 1 to 4, and then gradually stabilizes. This trend indicates that additional replicated examples yield diminishing returns in improving attack performance.
The impact of the feature extraction layer on attack performance is shown in Fig. 5. The fourth layer of the VGG-11 model demonstrates the most effective attack performance, while the first layer of the ResNet-18 model has the best attack effect. This difference is mainly due to the differing attention to feature information across models, which is related to their structure. Therefore, when using this method, the selection of the feature extraction layer should be selected to the specific structure of the surrogate model to achieve better attack performance.
Figure 5: (a) the number of iterations, (b) the number of copies, and (c) the impact of different model feature extraction layers on the attack success rate.
Figure 6: The impact of (a) the weight coefficient of feature loss and (b) the feature enlargement factor of the original examples on the attack success rate across different models.
@@
Figure 7: The impact of using different division parameters in the spatial and spectral dimensions on the attack success rate.
The impact of the weight coefficient of feature loss on attack performance is shown in Fig. 6. A weight coefficient of 0 indicates that no auxiliary information from the output layer is used, which is observed to degrade attack performance. When the weight coefficient is set to 0.03 or 0.05, the attack effect is the best. With 0.07 as the boundary, the weight of auxiliary loss increases and the effect gradually deteriorates. Therefore, setting the weight to 0.03 or 0.05 yields a more effective attack.
The impact of the original feature enlargement factor is shown in Fig. 6, with the enlargement factor ranging from 1 to 3. On the VGG-11 model, when the magnification is 2.7 or 3, the attack performance is the best; however, on the ResNet-18 model, the best performance occurs at a magnification of 1.2. This further shows that the effect of feature loss is affected by the model feature extraction method, indicating that optimal parameters can be selected according to the information of the surrogate model.
The influence of division parameters on attack performance is shown in Fig. 7. The range for spatial division parameters was set to 1-5, while that for spectral division was set to 1-4. As the spatial division number increased to 3, the attack performance improves significantly and then stabilized. Similarly, as the spectral division number increased to 3, the improvement in attack performance gradually diminished. To achieve a balance between computational efficiency and time cost, both the spatial and spectral division parameters were set to 3 in this work.
@@ V Conclusion
For the HSI classification problem, existing methods for generating adversarial examples often suffer from overfitting to the surrogate model and fail to fully utilize the features extracted during training. Building on this, we propose a comprehensive attack method to improve the transferability of adversarial examples in HSI classification tasks. Considering that HSIs have a large number of spectral bands, we divide the input hyperspectral examples into blocks along both spatial and spectral dimensions and apply various transformations to each block. Additionally, we design a weighted feature distancing loss function, which updates the perturbation by calculating the average gradient of multiple transformed images. It effectively suppresses the important features in the adversarial examples and enhances the transferability of adversarial attacks on HSIs. Experiments show that the adversarial examples produced by this method achieve significant transferability on multiple black-box models, which not only deepens the understanding of adversarial vulnerabilities in HSI classification but also provides an important reference for developing more robust defense strategies in this field.

### v2

@@ IV-D Ablation Experiment
SpatialT SpectralT WFDL-unEnlarge WFDL VGG-11 ResNet-18 VGG-19 Inc-V3 IncRes-V2
@@
SpatialT SpectralT WFDL-unEnlarge WFDL VGG-11 ResNet-18 VGG-19 Inc-V3 IncRes-V2
@@
Although performance slightly varies across different models, the attack effectiveness generally improves with the application of our proposed method. The experimental results show that partitioning the data only along the spectral dimension and applying transformations also reduces the model’s classification accuracy. By combining both spatial and spectral dimensions for 3D structure-invariant transformations, the transferability on the two datasets improved by approximately 4% and 6%, respectively. In addition, the weighted feature divergence loss, which takes into account the differences between feature channels, effectively improves the adversarial transferability. Furthermore, the operation of enlarging the original feature values when calculating the differences between feature maps further enhances the attack effectiveness of the proposed loss, thereby validating the effectiveness of the method.
IV-E Parameter Sensitivity Analysis Experiments
@@
To gain insight into the performance of the method, we further analyze the influence of various hyperparameters on the experimental results. The experiments were conducted mainly on the PaviaU dataset, with VGG-11 selected as the substitute model for generating adversarial examples, while the other models were treated as black-box models. The perturbation budget was set to 0.03.
The influence of division parameters on attack performance is shown in Fig. 6. The range for spatial division parameters was set to 1-5, while that for spectral division was set to 1-4. The blue line represents the impact of varying spatial division parameters on overall accuracy when the spectral dimension is not divided. Lines of other colors represent different spectral division parameters. Comparing the four lines across all models, the blue line exhibits significant fluctuation with parameter changes and consistently achieves the highest overall accuracy. The other lines are below, indicating that the operation of data chunking in the spectral dimension can improve the success rate of attacks. The distance between the lines reflects the enhanced attack effectiveness gained from spectral division. In the VGG-11 model, the green line (spectral division parameter 3) reaches its lowest point when the spatial division parameter is also 3, demonstrating the strongest attack effectiveness on the white-box model under this configuration. For other black-box models, most lines decrease as the spatial division parameter increases, this shows that the combination of the two has a good promoting effect. On some models, the lowest accuracy is achieved when the spectral division parameter is 4 and the spatial parameter is 4 or 5. It is observed that the results for spectral division parameters 3 and 4 (green and red lines) are quite similar. Considering that increased division inevitably raises computational complexity, and to balance attack effectiveness with computational efficiency, we sets both the spatial and spectral division parameters to 3.
@@
Figure 6: The impact of different division parameters used in the spatial and spectral dimensions on the results.
The impact of the feature maps extraction layer $l$ on attack performance is shown in Fig. 7. The fourth layer of the VGG-11 model demonstrates the most effective attack performance, with classification accuracy showing a declining trend as deeper layer feature maps are used. However, the first layer of the ResNet-18 model has the best attack effect. Notably, using feature maps from other layers in ResNet-18 has a relatively minor impact on classification accuracy. We analyze this discrepancy from the perspective of model architecture: the VGG-11 model utilizes simply stacked convolutional layers, whereas the ResNet-18 model incorporates residual connections (skip connections). Since the VGG-11 model lacks skip connections, the feature maps exhibit greater variation between different layers. Given that the results show that using deeper features generates adversarial examples with stronger transferability, it indicates that these features are more abstract and generalizable. Conversely, in ResNet-18 model, intermediate layer feature maps exhibit greater similarities due to the skip connections. Consequently, the choice of intermediate layer feature maps has a smaller effect on attack performance within ResNet-18. In this paper, based on experimental results across four corresponding layers, we selected the layer 3 as a compromise result for both substitute models to ensure relative fairness.
Fig. 7 shows the effect of iteration steps $I$ on attack performance. The iterations range from 1 to 60. When the number of iterations exceeds 20 times, the classification accuracy hardly decreases any more, and the improvement of attack effect is very limited. Therefore, to optimize computational costs while considering the attack’s efficacy, we set the number of iterations to 20.
The impact of the number of copies $N$ on attack performance is shown in Fig. 7. The number of transformed copies ranges from 1 to 20. As the number of transformed copies increases, the accuracy decreases most rapidly from 1 to 6, and then gradually stabilizes. This phenomenon indicates that additional transformed examples have little effect on improving attack performance. Therefore, in order to maintain a stable attack effect, we set N to 10.
Figure 7: The impact of (a) the network extraction layer parameter $l$, (b) the number of iterations and (c) the number of copies on the results.
The impact of the weight coefficient of cross-entropy classification loss $\eta$ on attack performance is shown in Fig. 8. When $\eta$ is set to 0, it indicates that the classification information of the output layer is not used. For nearly all models, classification accuracy is the highest at this point but subsequently shows a declining trend, indicating that appropriately incorporating the classification loss can enhance attack effectiveness. Whether using VGG-11 or ResNet-18 as the substitute model, the strongest attack effectiveness against black-box models is achieved when the weight $\eta$ is set to 0.01 or 0.03. With 0.07 as the boundary, as the weight of the cross-entropy classification loss increases beyond this value, the attack effectiveness against black-box models decreases. Therefore, setting $\eta$ to 0.03 yields superior experimental results.
The impact of the original feature values enlargement factor $\lambda$ is shown in Fig. 8, with the enlargement factor ranging from 1 to 3. When the VGG-11 model serves as the substitute model, the classification accuracy on the black-box model consistently exhibits a declining trend. The attack performance is best when $\lambda$ is 3. However, for the ResNet-18 model, optimal performance occurs at a magnification factor of 1.2. Enlarging feature values as a method to optimize the perturbation direction, although it has limited impact on improving transferability attacks, presents a new approach. The improvement effect of this method on different substitute models also has differences, demonstrating that the impact of weighted feature divergence loss on results is closely tied to the model’s architecture. In this paper, a small and valid value of 1.2 is selected, but the optimal parameter can also be selected according to the information of the substitute model.
Figure 8: The impact of (a) the weight coefficient $\eta$ of the cross-entropy classification loss and (b) the enlargement factor $\lambda$ of the original example features on the results.
IV-F Visualization Analysis
To gain clearer insight into the internal differences between adversarial examples generated by our method and those produced by other approaches, we extracted and compared the feature maps of the example from the VGG-11 model, as shown in Fig. 9. A darker orange indicates higher activation values, suggesting that these regions exert greater influence on the model’s classification performance. Although the other methods sometimes shift the locations of originally high-activation regions, they do not markedly suppress them. In contrast, although there is no obvious difference between the pseudo-color images of the two examples, the adversarial example feature map generated by our method shows the biggest difference with the original feature map: the regions most attended to in the clean example are largely replaced by lighter colors and thus become far less salient. This prevents the model from focusing on the most critical regions and thereby achieves the desired attack effect.
Original
$S$
SIA
BSR
Ours
Figure 9: The first row shows the false color images of the original example and its adversarial examples, while the second row shows the feature maps extracted from the VGG-11 model. The first column represents the original example, followed by the adversarial examples generated by $S$, SIA, BSR, and our method. Two sets of examples are shown here.
@@ V Conclusion
For the HSI classification problem, existing methods for generating adversarial examples often suffer from overfitting to the substitute model and fail to fully utilize the feature maps extracted during training. Building on this, we propose a comprehensive attack method to improve the transferability of adversarial examples in HSI classification tasks. Considering that HSI contains a large number of spectral bands, we perform 3D block partitioning on the input samples in both spatial and spectral dimensions, and apply random transformations to each block. Additionally, we design a weighted feature divergence loss that enhances the difference between the feature maps of the original and adversarial examples, suppressing features related to the true class in the adversarial example, thereby improving transferability. Experiments show that the adversarial examples produced by this method achieve significant transferability on multiple black-box models, which not only deepens the understanding of adversarial vulnerabilities in HSI classification but also provides an important reference for developing more robust defense strategies in this field.